Can you please answer me this algebra question
Roy is 11 years old and his uncle is 59 years old. How many years ago was Roy's uncle 7 times old as Roy?

Answer by ewatrrr(24785)

x How many years ago
59-x = 7(11-x)
59-x = 77 - 7x
6x = 18
x = 3, How many years ago Roy's uncle was 7 times old as Roy
...
And...checking ...3 years ago
56 = 7(8)